### v1

@@ 3.4 Regularizing Expert Routing
where $\operatorname{stopgrad}(\cdot)$ is the stop-gradient operator to prevent error propagation of this loss term when the coefficient of variation is less than a pre-defined threshold $\gamma=0.1$. While this loss was initially proposed for balancing computational budgets [38, 36], we adapt its use for promoting expert specialization. We add an additional threshold constraint due to our instance-wise routing, which is more likely to have a larger coefficient of variation than per-token routing.
4 Experiments
4.1 Implementation Details
@@
Architecture Details. For all experiments, unless otherwise specified, we use Swin-B [28] as the vision encoder and Bert-base [3] as the text encoder. Vision is used as main modality. Following the experiment setup in [11, 21], we also finetune a linear head for each dataset. We implement the mapper $f_{m}(\cdot)$ as a two-layer MLP with GeLU [9] nonlinearily. Regarding prompt tuning, we use $l=6$ prompts and $k=16$ experts by default, and the tunable prompt is applied to all layers of the main modality encoder. Vanilla prompt tuning [20] with $l^{\prime}=4$ is used to tune the complementary modality.
Training Details. All images are resized and cropped to the size of $224\times 224$. All models are trained for 15 epoches, using the AdamW [29] optimizer with a learning rate of $4\times 10^{-4}$ for vision and $5\times 10^{-4}$ for text. We use a constant step decay scheduler that scales the learning rate at epochs 3 and 6 by a factor of $\gamma=0.4$. All models are trained on a server with a RTX4090 GPU. The mean and standard deviation of metrics, measured over 4 different experiments are reported when possible.
4.2 Datasets
MM-IMDB [1] is a multimodal movie classification dataset. It consists of 25,956 image-text pairs, where each pair includes a movie poster and a plot summary The dataset supports multi-label genre classification across a spectrum of 23 genres with a long-tail distribution.
UPMC Food-101 [41] is a comprehensive multimodal dataset designed for fine-grained recipe classification with 90,840 image-text pairs for 101 food classes.
SNLI-VE [43] is a large-scale multimodal dataset with 565,286 image-text pairs. The task for this dataset is visual entailment, which requires the model to infer whether a hypothesis matches the given premise. In our experiments, following [21], we only use the image premise. However, it is worth noting that some other works may also include the text premise.
4.3 Compared Methods
First, we compare our method with several baselines for multimodal fusion.
@@ 4.3 Compared Methods
ImgOnly / TextOnly. Only fine-tune one encoder, and the input of the other modality is discarded.
P-ImgOnly / P-TextOnly. Only prompt-tune one encoder, and the input of the other modality is discarded.
@@
P-LateConcat. Similar as LateConcat but prompt-tune each encoder instead of fine-tuning. However, the linear head is still fine-tuned.
SequentialFuse. This method first extracts features from the complementary modality and maps them to the embedding space of the main modality encoder in the same way as our mapped prompt for end-to-end training. Both encoders are fine-tuned. This is a strong baseline that can be considered as our method without MoPE, but with all parameters fine-tuned.
P-SequentialFuse. Prompt-tune each encoders in SequentialFuse.
For a fair comparison, all the prompt-tuning methods listed above use the same prompt length ($l=6$) as our method. In addition to these baselines, we also compare our methods with existing prompt-based fusion methods, including MMBT [15], Frozen [39], PromptFuse and BlindPrompt [22], and PMF [21]. Among these methods, the setting in PromptFuse [22] is the most similar to ours, as it also assumes one modality to be a black-box.
Table 1: Quantitative results on three multimodal dataset. We compare all baseline methods, existing prompt-based fusion methods, and our method with different expert numbers. Our method achieve state-of-the-art performance and parameter-efficiency. (†): Our implementation with prefix-tuning.
Method Param SNLI-VE($\uparrow$) Food-101($\uparrow$) MM-IMDB($\uparrow$)
finetuning ImgOnly 86.9M 33.34$\pm$.03 75.64$\pm$.21 39.21$\pm$.32/53.85$\pm$.21
TextOnly 109.0M 69.58$\pm$.24 86.92$\pm$.21 58.80$\pm$.31/65.37$\pm$.13
LateConcat 196.0M 72.01$\pm$.23 93.19$\pm$.21 60.43$\pm$.68/67.77$\pm$.36
SequentialFuse 197.0M 74.28$\pm$.46 93.73$\pm$.16 59.22$\pm$.38/66.34$\pm$.19
MMBT [15] 196.5M 67.58 94.10 60.80/66.10
prompt-tuning P-ImgOnly 0.1M 33.34$\pm$.01 76.65$\pm$.07 33.70$\pm$.55/50.04$\pm$.27
P-TextOnly 0.1M 64.86$\pm$.32 81.01$\pm$.50 52.19$\pm$.32/61.16$\pm$.43
P-LateConcat 1.3M 64.29$\pm$.28 90.27$\pm$.40 56.95$\pm$.37/64.23$\pm$.26
P-SequentialFuse 1.1M 65.01$\pm$.18 81.27$\pm$.22 55.57$\pm$.63/63.98$\pm$.35
P-MMBT [21] 0.9M 67.58 81.07 52.95/59.30
PromptFuse [22] <0.1M 64.53 82.21 48.59/54.49
BlindPrompt [22] <0.1M 65.54 84.56 50.18/56.46
PromptFuse(†) [22] 0.1M 64.94 82.14 50.78/60.96
PMF [21] 2.5M 71.92 91.51 58.77/64.51
Ours (ViT-b, $k=4$) 1.6M 73.47$\pm$.11 91.55$\pm$.12 62.37$\pm$.35/68.73$\pm$.24
Ours (Swin-b, $k=4$) 1.6M 73.14±.$\pm.21$ 91.54$\pm$.21 61.93$\pm$.37/68.19$\pm$.14
Ours (Swin-b, $k=16$) 2.6M 73.59$\pm$.15 92.05$\pm$.11 62.01$\pm$.21/68.24$\pm$.12
Table 2: Ablation on three types of prompt disentanglement. The results from all combinations of mapped prompt, dynamic prompt, and static prompt are presented. Our full method with all prompts achieves the best result.
Prompt SNLI-VE($\uparrow$) Food-101($\uparrow$) MM-IMDB($\uparrow$)
@@
$[\mathbf{P}_{s},\mathbf{P}_{d},P_{m}]$ $\mathbf{73.59\!\pm\!.15}$ $\mathbf{92.05\!\pm\!.11}$ $\mathbf{62.01\!\pm\!.21}$/$\mathbf{68.24\!\pm\!.12}$
4.4 Main Results
The quantitative results of all baselines, compared methods, and our method with different expert numbers $k$ are summarized in Tab 1. The metric on SNLI-VE and Food-101 is accuracy (%), and MM-IMDB is F1-Macro and F1-Micro (%). We also list the number of trainable parameters (million) for each method.
Our method outperforms all prompt-based methods, and is competitive with fine-tuning. Specifically, when compared with the fine-tuning baselines, SequentialFuse and LateConcat, our method delivers competitive accuracy on the Food-101 dataset and superior results on the SNLI-VE and MM-IMDB datasets, while requiring as few as 0.8% of the trainable parameters.
Apart from surpassing baselines, the proposed method also outperforms existing prompt-based fusion methods, including PromptFuse [22], BlindPrompt [22], and PMF [21]. Notably, our method outperforms the current SOTA, PMF [21], by a significant margin on all datasets, while requiring $37\%$ fewer parameters. Moreover, PMF assumes both encoders share the same Transformer architecture. By contrast, our method has weaker assumptions, which could be easily extended to any Transfomer-based architecture, such as ViT [4] and Swin [28].
4.5 Qualitative Result of MoPE Routing
During end-to-end tuning, prompt experts in MoPE spontaneously specialize in different concepts. Fig. 3 presents two examples of this specialization. Specifically, we visualize the expert with the highest score as the routed expert for each instance pair. In the provided examples, we observe that expert-4 specialize in children-related concepts, while expert-12 focuses on crowds. These examples also demonstrate the interpretability of our MoPE-based prompting, as opposed to black-box conditional prompt such as CoOp [50] and CoCoOp [49]. We encourage readers for a more comprehensive visualizations in Section 0.C.2.
5 Analysis and Discussion
@@
Our three-types of disentangled prompts are collaborative. We first ablate each type of prompt in our prompt disentanglement, and the results are summarized in Tab. 2. Our full method achieves the best performance, indicating that the prompts collaborate effectively with each other. Specifically, the dynamic prompt is essential for multimodal fusion as it learns the interplay between different modalities. This is in contrast with previous methods [22, 39, 45], where each modality do not explicitly interact. Adding the dynamic prompt induces a significant gain of 13.5%, 13.26, 8.4%/6.5% on all datasets. Without the mapped prompt, the fine-grained information from complimentary modality is lost, leading to a significant drop in all datasets. The static prompt is also necessary, serving as a special expert that is always routed. Adding this special expert allows the other experts to focus more on capturing instance-level features, resulting in a slight performance gain.
@@
Expert-scaling is more expressive than length-scaling. To study the expressiveness of MoPE, we compare expert-scaling with length-scaling. Our starting point for MoPE is $l=6$ prompts and $k=2$ experts (when $k\leq 1$, MoPE degenerate into vanilla prompt), which account for $(2+1)\times 6+1=19$ tunable prompts in total per-layer. We increase the number of $k$, and at each step we also report the result of simply increasing $l$ in vanilla prompt, up to the same total number of prompts. The results are summarised in Fig. 4.
@@
As illustrated in the figure, adding the MoPE design with as few as $k=2$ experts boosts performance, and increasing the number of experts leads to a monotonic performance improvement. Although our method may also saturate when the number of experts becomes excessive, it does not face the over-length deterioration issue observed in vanilla prompt tuning. In contrast, merely augmenting the number of prompts $l$ in vanilla prompt tuning does not lead to a linear performance improvement and is prone to performance deterioration. This observation is consistent with previous findings [11, 44, 14, 20]. This makes the prompt length $l$ a hyperparameter that is difficult to tune manually. In addition to accuracy considerations, longer prompts in vanilla prompt tuning can exacerbate computational overhead due to the quadratic time complexity of self-attention. This is why we cannot experiment with $l>199$ for the other two methods. On the other hand, MoPE scales by conditioning the dynamic prompt on more experts while maintaining a fixed length of learnable prompts during self-attention. As a result, our method maintains a nearly constant time complexity.
Table 3: Result of dense routing v.s. sparse routing. Dense routing achieves slightly better results on all three datasets. Routing SNLI-VE $\uparrow$ Food-101 $\uparrow$ MM-IMDB $\uparrow$ Dense 73.55 91.74 62.01/68.25 Sparse 71.15 90.89 60.28/66.01 Table 4: Different way of generating routing embedding. Frozen routing embedding with orthogonal initialization is as good as learned one. Routing Embed SNLI-VE $\uparrow$ Food-101 $\uparrow$ MM-IMDB $\uparrow$ Frozen 73.55 91.74 62.01/68.25 Learned 73.13 91.20 61.64/67.97
@@
[Uncaptioned image] Figure 5: Scaling performance with increased training data. We show the performance of our method and representative methods as we progressively increase the amount of training data, or “#shots”. Our method outperform other prompt-tuning methods at all data scales.
@@
MoPE scales better when more training data is available. Previous prompting methods have been shown to not scale well with respect to increased training data [7, 22, 39, 44]. We study how the proposed MoPE method performs on different data scales to assess its scalability. Specifically, we sub-sample the training-set with different shots to represent a range of low-shot to high-shot learning scenarios. We then train our method and other prompt-tuning and fine-tuning methods with the same subsampled data. The mean accuracy and F1-Micro score on SNLI-VE and MM-IMDB, respectively, are reported in Fig. 5.
@@
As the figure shows, our MoPE-based method is the most scalable, achieving superior results across all data scales. Specifically, we consistently match the results of the fine-tuning method, SequentialFuse, on the SNLI-VE dataset, and we achieve superior results on the MM-IMDB dataset. By contrast, the method based on vanilla-prompt does not scale well with respect to increased training data, and there is a consistent performance gap between ours and the compared method. This performance gap becomes more significant on larger datasets, such as when training with $10^{5}$ shots or using the full training set.
Refer to caption Figure 6: Routing determinism. We show the entropy of routing score during training. Dense routing are more deterministic as optimization progress. Better viewed with color.
@@
Dense routing with frozen routing embedding achieves the best result. Previous MoE with FFN acting as experts usually employ a sparse gating function that only select one or a few experts per token, while we use a dense router (Eqn. 4) that weigh all experts. To compare the effect, we follow  [36] to implement a sparse router by adding a Top-1 gate: $\mathbf{r}^{\prime}=\operatorname{TOP}_{1}(\mathbf{r})$.
The results of using sparse routing versus dense routing are reported in Tab. 5. Our experiments indicate that using the dense router yields marginally better results.
To further understand how dense routing benefits the model, we evaluated the degree of uncertainty in routing by calculating the empirical entropy of the routing score $\mathcal{H}(\mathbf{r})=-\sum_{i=1}^{n}r_{i}log_{2}(r_{i})$, averaged across training batches and all layers. In essence, high entropy indicates that routing behaves more randomly. The relationship between entropy and optimization steps is illustrated in Fig. 6. As demonstrated in the figure, dense routing is more deterministic.
@@
In addition to routing sparsity, we also investigate the effect of using frozen routing embeddings. As Tab. 5 shows, fixed routing embedding could achieve the same result as the trained one.
@@
Importance loss aids expert specialization. The importance loss is crucial for avoiding degenerate routing solutions and, consequently, aids expert specialization. In Fig. 7, we visualize how the importance of each expert (i.e., average routing score) changes when training on the SNLI-VE dataset. Without the importance loss, routing adheres to its initial state, resulting in a skewed distribution where a few experts dominate. The importance loss allows other experts to be used by penalizing highly unbalanced expert importance.
Architecture Tuning MM-IMDB ($\uparrow$)↑ )
BoW* Fine-tune 48.20/57.50
Bert [3] Frozen 58.86/66.13
GPT-2 [35] Frozen 34.03/50.84
Longformer [3] Bitfit [46] 60.34/67.27
Table 5: Our fusing method is highly modular. We treat complementary modality as black-box, allowing different models to be plug-in for fusion. (*): Bag-of-word initialized with Bert word embeddings.
Our fusion method are highly modular. Our method abstracts the complimentary modality as a black-box, allowing high modality of both modalities. In particular, our method allow arbitrary models to be seamless plug-in for multimodal fusion. We give some results in Tab. 5.
@@
How could MoPE be more expressive & Limitations. We have demonstrated that the MoPE-based method is more expressive for multimodal fusion. Compared with existing prompt-based methods, our approach scales better with data and trainable parameters. This increased expressiveness comes from two factors. First, the prompts used in previous methods are not adaptive, meaning that a single prompt is used for all instances. Finding a instance-wise optimal prompt could lead to better results. Secondly, we scale the prompt by using more experts, since a long prompt can be difficult to optimize [32] for optimal performance. By doing so, we shift the difficulty from learning a universal long prompt to learning a parameterized router with multiple short prompts. As a result, our empirical expressiveness is closer to the theoretical upper bound of prompt-tuning.
Despite improving empirical expressiveness, our prompt still functions in the same way as vanilla prompts during attention, resulting in same theoretical upper-bound of expressiveness. This is also reflected by the small gap between ours and fine-tuning in large datasets such as SNLI-VE and UPMC Food-101. Future studies could focus on designing a more expressive prompting technique.
6 Conclusion

### v2

@@ Introduction
where $\operatorname{\sigma}(\cdot)$ is the stop-gradient operator to prevent error propagation of this loss term when the coefficient of variation is less than a pre-defined threshold $\gamma=0.1$. While this loss was initially proposed for balancing computational budgets (Shazeer et al. 2017; Riquelme et al. 2021), we adapt its use for promoting expert specialization. We add an additional threshold constraint due to our instance-wise routing, which is more likely to have a larger coefficient of variation than per-token routing.
I+L A+L A+V+L
Method Param Speed (ms) SNLI-VE UPMC Food MM-IMDB MUsTARD MUsTARD
Acc ($\uparrow$) Acc ($\uparrow$) F1-macro/F1-micro ($\uparrow$) Pre ($\uparrow$) Pre ($\uparrow$)
fine-tuning ImgOnly 86.9M 16.45$\pm$0.2 33.34 75.64 39.21/53.85 - -
TextOnly 109.0M 7.28$\pm$0.9 69.58 86.92 58.80/65.37 65.41 65.41
LateConcat 196.0M 22.79$\pm$0.9 72.01 93.19 60.43/67.77 68.82 69.40
SeqFuse 197.0M 23.32$\pm$1.2 74.28 93.73 59.22/66.34 68.13 71.35
MMBT 196.5M 15.91$\pm$1.2 67.58 94.10 60.80/66.10 - -
prompt-tuning P-ImgOnly 0.1M 21.01$\pm$1.1 33.34 76.65 33.70/50.04 - -
P-TextOnly 0.1M 7.57$\pm$1.2 64.86 81.01 52.19/61.16 59.27 59.27
P-LateConcat 1.3M 28.59$\pm$2.1 64.29 90.27 56.95/64.23 60.71 65.17
P-SeqFuse 1.1M 28.52$\pm$1.9 65.01 81.27 55.57/63.98 63.72 65.73
P-MMBT 0.9M 16.71$\pm$1.3 67.58 81.07 52.95/59.30 - -
PromptFuse <0.1M 28.75$\pm$0.9 64.53 82.21 48.59/54.49 63.76 64.20
BlindPrompt <0.1M 29.52$\pm$1.2 65.54 84.56 50.18/56.46 62.01 63.80
PromptFuse(†) 0.1M 29.26$\pm$1.7 64.94 82.14 50.78/60.96 64.73 66.29
PMF 2.5M 32.21$\pm$0.8 71.92 91.51 58.77/64.51 - -
Ours ($k\!=\!4$) 1.6M 30.47$\pm$1.4 73.14 91.54 61.93/68.19 67.12 67.35
Ours ($k\!=\!16$) 2.6M 30.44$\pm$1.3 73.59 92.05 62.01/68.24 68.73 69.94
Table 1: Quantitative results on multimodal classification. Our method achieve the best performance and parameter-efficiency against all prompt-fusion methods. I: Image, L: Language, A: Audio, V: Video. The main modality is underlined, and the best prompt tuning method is in bold, and (†): Our re-implementation with prompt applied to all layers.
Method Param RefCOCO, mIOU($\uparrow$) RefCOCO+, mIOU($\uparrow$)
val testA testB val testA testB
SeqFuse 231.0M 53.48 55.76 52.03 40.22 42.2 37.91
P-SeqFuse 35.1M 47.69 46.23 45.81 30.66 31.48 28.79
PromptFuse(†) 35.1M 43.23 39.71 47.74 27.72 33.37 23.67
Ours 35.5M 58.40 60.03 53.23 43.80 46.12 38.88
Table 2: Quantitative result on referring image segmentation We report the total number of parameters in million (including the trainable segmentation head), and metrics (mIOU) on RefCOCO and RefCOCO+.
Experiments
Datasets. Following previous prompt fusion methods (Liang, Zhao, and Schütze 2022; Li et al. 2023), we conduct experiments on MM-IMDB (Arevalo et al. 2017) and UPMC Food-101 (Wang et al. 2015) for classification, SNLI-VE (Xie et al. 2019) for visual entailment (with image premise only), MUsTARD (Castro et al. 2019) for sarcasm detection in speaker-dependent setting. We additionally test our method on the RefCOCO and RefCOCO+ (Kazemzadeh et al. 2014) dataset for referring expression segmentation.
@@
Architecture Details. In main experiment, we use Swin-B (Liu et al. 2021b) as the image encoder, Bert-base (Devlin et al. 2018) as the text encoder, Wav2Vec2 (Baevski et al. 2020) as the audio encoder. For the video modality, we use the same frame averaging technique by adapting a ViT (Dosovitskiy et al. 2020) encoder, which follows PromptFuse (Liang, Zhao, and Schütze 2022). Image is used as main modality unless otherwise specified. Following the experiment setup in (Jia et al. 2022; Li et al. 2023; Liang, Zhao, and Schütze 2022), we finetune dataset-specific heads. We implement the mapper as a two-layer MLP with GeLU nonlinearily. Regarding prompt, we set $l=6$ and use $k=16$ experts by default, and the prompts are applied to all layers of the main modality encoder. Vanilla prompt tuning (Li and Liang 2021) with $l^{\prime}=4$ is used to tune the $\mathcal{E}_{\mathbb{Y}}$.
Training Details. All models are trained for 20 epoches, using the AdamW (Loshchilov and Hutter 2017) optimizer with a learning rate of $4\times 10^{-4}$ for main modality and $5\times 10^{-4}$ for complementary modality. All models are trained with a RTX-4090 GPU.
Further experiment details could be found in the Supplementary Material.
Compared Methods. We consider following baselines:
@@
ImgOnly / TextOnly. Fine-tune one encoder only, and the input of the other modality is discarded.
P-ImgOnly / P-TextOnly. Only prompt-tune one encoder.
@@
P-LateConcat. Similar as LateConcat but prompt-tune each encoder instead of fine-tuning.
SeqFuse. This baseline first extracts features from the complementary modality and maps them to the embedding space of the main modality encoder by a MLP. Both encoders are fine-tuned. This is a strong baseline that can be considered as our method without MoPE, but with all parameters fine-tuned.
P-SeqFuse. Prompt-tune each encoders in SeqFuse.
In addition to these baselines, we also compare our methods with existing prompt-based fusion methods, including MMBT (Kiela et al. 2019), Frozen (Tsimpoukelli et al. 2021), PromptFuse and BlindPrompt (Liang, Zhao, and Schütze 2022), and PMF (Li et al. 2023). Among these methods, the setting in PromptFuse (Liang, Zhao, and Schütze 2022) is the most similar to ours, as it also assumes one modality to be a black-box and is by design compatible with more than two modalities. To ensure a fair comparison, all prompt-tuning baselines utilize the same prompt length ($l=6$) as our method.
Prompt SNLI-VE($\uparrow$) UPMC Food($\uparrow$) MM-IMDB($\uparrow$)
@@
$[\mathbf{P}_{s},\mathbf{P}_{d},P_{m}]$ $\mathbf{73.59}{\color[rgb]{.5,.5,.5}{\pm.15}}$15 $\mathbf{92.05}{\color[rgb]{.5,.5,.5}{\pm.11}}$11 $12\mathbf{62.01}{\color[rgb]{.5,.5,.5}{\pm.21}}/\mathbf{68.24}{\color[rgb]{.5,.5,.5}{\pm.12}}$12
Table 3: Ablation on prompt decomposition. The results from all combinations of mapped prompt, dynamic prompt, and static prompt are presented. Our full method with all prompts achieves the best result.
Quantitative Comparison with Previous Methods
Multimodal Classification. The quantitative results on multimodal classification on all methods are summarized in Tab 1. The metric on SNLI-VE, UPMC Food-101 is accuracy (%), MM-IMDB is F1-Macro and F1-Micro (%), and on MuSTARD is precision (%). We also list the inference time (in milliseconds) and number of trainable parameters (in million) for each method.
Our method outperforms all prompt-based methods, and is competitive with fine-tuning. Specifically, when compared with the fine-tuning baselines, SeqFuse and LateConcat, our method delivers competitive accuracy on the UPMC Food-101 dataset and superior results on the SNLI-VE, MM-IMDB and MusTARD datasets, while requiring as few as 0.8% of the trainable parameters.
The proposed method also outperforms all existing prompt fusion methods, including PromptFuse (Liang, Zhao, and Schütze 2022), BlindPrompt (Liang, Zhao, and Schütze 2022), and PMF (Li et al. 2023). Notably, our method ($k=4$) outperforms the current SOTA, PMF, by a significant margin on all datasets, while requiring $37\%$ fewer parameters. Moreover, our method has a weak assumption than PMF, which assumes both encoders share the same Transformer architecture, making it hard to be extended to other modalities and other model architectures.
Referring Expression Segmentation. Previous prompt fusion methods mainly test their performance on classification (Li et al. 2023; Liang, Zhao, and Schütze 2022; Tsimpoukelli et al. 2021). The improved expressiveness of MoPE allows learning of harder tasks, for example, referring image segmentation. Following experiment setting in VPT, We additionally finetune a SeTR-PUP (Zheng et al. 2021) head for segmentation. The result is summarized in Tab. 2. Our proposed method achieves the best result among all compared methods. In particular we outperform the fine-tuning baseline SeqFuse, which suffers from catastrophic forgetting problem on this task.
Qualitative Result of MoPE Routing
Through training, prompt experts in MoPE spontaneously specialize, and we present two examples in Fig. 3. We visualize the expert with the highest score as the routed expert for each instance pair. In the provided examples, we observe that expert-4 specialize in children-related concepts, while expert-12 focuses on crowds. These examples also demonstrate the interpretability of MoPE, as opposed to black-box conditional prompt such as CoOp (Zhou et al. 2022b). More visualizations could be found in supplementary material.
Analysis and Discussion
@@
Our decomposed prompts are collaborative. We ablate each type of prompt in our instance-wise prompt decomposition, with results across 3 random seeds reported in Tab. 3 . Our full method achieves the best performance, indicating effective prompt collaboration. Adding the dynamic prompt yields gains of 13.5%, 13.26%, 8.4%/6.5% on all datasets, this is different from previous methods (Liang, Zhao, and Schütze 2022; Tsimpoukelli et al. 2021) where modalities don’t explicitly interact. Without the mapped prompt, fine-grained information from complementary modalities is lost, causing significant drops. The static prompt is also necessary, acting as a special expert always routed for capturing global-level features.
@@
Expert-scaling is more expressive than length-scaling. To study the expressiveness of MoPE, we compare expert-scaling with length-scaling. Our starting point for MoPE is $l=6$ prompts and $k=2$ experts, which account for $(2+1)\times 6+1=19$ tunable prompts in total per-layer. We increase the number of $k$, and at each step we also report the result of increasing $l$ in vanilla prompt to the same total number of prompts. The results are summarized in Fig. 4.
@@
As illustrated in the figure, adding the MoPE design with as few as $k=2$ experts boosts performance, and increasing the number of experts leads to a monotonic performance improvement which avoids over-length deterioration. In contrast, length-scaling does not lead to a linear performance improvement and is prone to performance deterioration, which is consistent with previous findings (Jia et al. 2022; Yang et al. 2022a; Khattak et al. 2023; Li and Liang 2021). Additionally, longer prompts in vanilla prompt tuning can exacerbate computational overhead due to the quadratic time complexity of self-attention. This is why we cannot experiment with $l>199$ for the other two methods. By contrast, MoPE scales by conditioning the dynamic prompt on more experts while maintaining a fixed prompt length during actual self-attention. As a result, our method maintains a nearly constant time complexity.
@@
Figure 5: Scaling performance with increased training data. We show the performance of our method and representative methods as we progressively increase the amount of training data, or “#shots”. Our method outperform other prompt fusion methods at all data scales.
@@
MoPE scales better with more training data. Previous prompting methods have been shown to not scale well with respect to increased training data (Gao et al. 2022; Liang, Zhao, and Schütze 2022; Tsimpoukelli et al. 2021; Yang et al. 2022a). To assess the scalability of MoPE, we sub-sample the training set with different shots to simulate a range of low-shot to high-shot learning scenarios. We then train our method and other prompt-tuning and fine-tuning methods on the same subset of SNLI-VE and MM-IMDB.
@@
As Fig. 5 shows, our MoPE-based method demonstrates superior scalability compared to other prompt tuning methods. Specifically, we consistently match the results of the fine-tuning method, SeqFuse, on the SNLI-VE dataset, while surpassing it on the MM-IMDB dataset. By contrast, the methods based on vanilla prompts do not scale well with respect to increased training data, resulting in a consistent performance gap between our method and the compared methods. This performance disparity becomes more pronounced on larger datasets, such as when training with $10^{5}$ shots or using the full training set.
@@
Frozen routing embedding outperform learned embedding. We ablate the effect of using frozen routing embeddings. As Tab. 4 shows, fixed routing embedding could achieve the same result as the trained one.
Routing Embed SNLI-VE $\uparrow$ UPMC Food $\uparrow$ MM-IMDB $\uparrow$
Frozen 73.55 91.74 62.01/68.25
Learned 73.13 91.20 61.64/67.97
Table 4: Result of frozen routing embedding. Frozen routing embedding are slightly better than learned one.
@@
Importance loss aids expert specialization. The importance loss is crucial for avoiding degenerate routing solutions and, consequently, aids expert specialization. In Fig. 6, we visualize how the importance (i.e., average routing score) of each expert changes when training on the SNLI-VE dataset. Without the importance loss, routing adheres to its initial state, resulting in a skewed distribution where a few experts are always being routed. Adding it aids expert specialization by penalizing highly unbalanced expert importance.
@@
How MoPE could be more expressive. We have demonstrated that the proposed MoPE scales better with trainable parameters and data. This increased expressiveness comes from two factors. First, previous methods use a shared prompt for all instances and neglect multimodal interplay. By contrast, our dynamic prompt with MoPE is adaptive, and finding an instance-wise optimal prompt can lead to better results. Secondly, MoPE circumvents the challenge of optimizing a long prompt (Petrov, Torr, and Bibi 2023). By doing so, we shift the difficulty from learning a universal long prompt to learning a parameterized router with multiple short prompts. As a result, our empirical expressiveness is closer to the theoretical upper bound of prompt tuning.
Conclusion